OpenStudy (anonymous):

What is the value of x over y when x=4/9 and y=5/3

OpenStudy (agent47):

x/y=4/9/5/3=(4/9)*(3/5)

OpenStudy (lgbasallote):

\[\implies \frac{\frac 49}{\frac 53} \implies \frac 49 \times \frac 35\] does that give you any hints?

OpenStudy (agent47):

\[{4 \over 9} \div {5 \over 3}\]now a/b / c/d = a/b * d/c

OpenStudy (agent47):

4/15